OpenStudy (anonymous):

Given that log(7) = 0.8451, find the value of the logarithm: log (1/7000)

OpenStudy (anonymous):

log(1/7000) = log1 - log7000 = -log(7*10^3) = -log7 + log10^3 = -log7 + 3 = 0.8451 + 3

OpenStudy (campbell_st):

there is a slight error in the calculation it should be - (log 7 + log 1000) =-(0.8451 + 3) = -3.8451

OpenStudy (anonymous):

yeah... thanks campbell_st :)

OpenStudy (anonymous):

that's what I got, and the stupid online assignment says its wrong

OpenStudy (anonymous):

computer, eh? can't trust 'em!!!

OpenStudy (campbell_st):

lol

OpenStudy (anonymous):

seriously...

OpenStudy (campbell_st):

the only question I have then... is it base 10 logs..?

OpenStudy (anonymous):

Have you tried using more decimal numbers? Like -3.84509. That shouldn't make a lot of sense, since they gave the answer already rounded up, but it does not hurt to give it a shot.

OpenStudy (anonymous):

yeah I tried different variations of the decimal answer.. didn't take it. Base 10 logs, how would that be shown?

OpenStudy (campbell_st):

what about leaving it as an exact answer -3 -log(7)

OpenStudy (anonymous):

hmmm.. nope, it can't compute the answer with "log" in it so I'm guessing its gotta be a decimal answer. this is lame. guess i can email the professor. I don't know what else she wants from me lol